Suppose that in 2020, the government runs a primary budget deficit of 2% of GDP, the nominal interest rate is equal to 5% per year and GDP grows at 3% per year.

(a) Starting with the government budget equation, derive the steady-state debt to GDP ratio.

(b) Is this steady state stable or unstable? Explain your answers.

(c) Explain the difference between the primary deficit and the reported deficit.

(d) Assume that in 2020, the government had a debt to GDP ratio of 25%. How big was the reported deficit to GDP ratio?


1
Expert's answer
2021-03-23T11:55:56-0400

a: Government budget function=difference between revenues (R) and expenditures(G).

=R-G

Primary budget deficit= government budget function.

Steady state debt ratio=rate of GDP growth *stock of debt(primary budget deficit*nominal interest rate)

=3*2*5=30%

b: unstable steady state .This is because a small disturbance like 2%primary budget deficit moves away GDP from equilibrium.

c:Reported deficit shows the rate at which the stock of debts is growing while primary deficit shows amount at which government expenditure exceeds tax revenue.

d: 30-25=5%